# Conformal Alignment: Knowing When to Trust Foundation Models with Guarantees

### v1

@@ Abstract
Before deploying outputs from foundation models in high-stakes tasks, it is imperative to ensure that they align with human values. For instance, in radiology report generation, reports generated by a vision-language model must align with human evaluations before their use in medical decision-making. This paper presents Conformal Alignment, a general framework for identifying units whose outputs meet a user-specified alignment criterion. It is guaranteed that on average, a prescribed fraction of selected units indeed meet the alignment criterion, regardless of the foundation model or the data distribution. Given any pre-trained model and new units with model-generated outputs, Conformal Alignment leverages a set of reference data with ground-truth alignment status to train an alignment predictor. It then selects new units whose predicted alignment scores surpass a data-dependent threshold, certifying their corresponding outputs as trustworthy. Through applications to question answering and radiology report generation, we demonstrate that our method is able to accurately identify units with trustworthy outputs via lightweight training over a moderate amount of reference data. En route, we investigate the informativeness of various features in alignment prediction and combine them with standard models to construct the alignment predictor.
@@ 1 Introduction
Towards uncertainty quantification of black-box models, conformal prediction [47, CP] offers a versatile, distribution-free solution. While CP has been applied to classification and regression tasks addressed by foundation models [23, 25, 43, 45, 51], its use to ensure the alignment of general-form outputs remains largely unexplored. Recently, [33] propose to generate multiple outputs for one input until it is guaranteed that at least one output meets a specific alignment criterion. Such a guarantee, however, may not be immediately practical, as users must still determine (manually) which output(s) is aligned. [29] modify the machine-generated outputs to ensure factuality with high probability, which in turn may sacrifice model power by making responses vague. As such, determining guarantees that are practically meaningful for downstream tasks and developing methods to achieve them remain an active research area.
This paper introduces a distinct type of guarantee: we aim to certify whether each model output is aligned or not, such that those certified are ensured to be mostly correct. This guarantee is directly relevant to downstream tasks, as it allows immediate and reliable use of the certified outputs without further modifications. On the other hand, we abstain from deploying model outputs for units that are not selected—intuitively, those are where the model lacks confidence (i.e., “knowing they do not know”), and they can be evaluated by human experts. Such a practice therefore provides an automated pipeline for safe deployment of model-generated outputs.
We present Conformal Alignment (CA), a general framework that determines when to trust foundation model outputs with finite-sample, distribution-free guarantees. The framework is based on CP ideas, but unlike standard CP that searches the sampling space to construct a prediction set that contains a desired output, we leverage the quantified confidence as an instrument to select units whose already-generated outputs are trusted. Given any model and any alignment criterion, Conformal Alignment ensures that a prescribed proportion of the selected units’ model outputs indeed meet the alignment criterion.
Concretely, suppose $m\in\mathbb{N}_{+}$ new units awaiting outputs. Given an error level $\alpha\in(0,1)$, Conformal Alignment controls the false discovery rate (FDR) in selecting units with trustworthy outputs:
@@
We illustrate the pipeline of Conformal Alignment via a use case in radiology report generation (detailed in Section 5) in Figure 1: each prompt corresponds to an X-ray scan and a vision-language model generates a report for each scan. Our framework identifies a subset of generated reports that are reliable in the sense that at least, say $99\%$, of the selected reports are guaranteed to be aligned if they were to be compared with a human expert report.
@@ 2 Problem setup
For a test set $\mathcal{D}_{\textnormal{test}}=\{X_{n+j}\}_{j\in[m]}$, we aim to select a subset ${\mathcal{S}}\subseteq[m]:=\{1,\dots,m\}$ such that most of their (unobserved) alignment scores are above a pre-fixed threshold $c\in\mathbb{R}$. This is measured by
@@
$\text{Power}=\mathbb{E}\Bigg{[}\frac{\sum_{j\in[m]}\operatorname{\mathds{1}}\{A_{n+j}>c,j\in{\mathcal{S}}\}}{\max(\sum_{j\in[m]}\operatorname{\mathds{1}}\{A_{n+j}>c\},1)}\Bigg{]}.$
It should be emphasized that our framework prioritizes FDR control over power, where we strictly enforce the former while optimizing the latter under the constraints. This is related to but slightly differs from the selection with controlled risk setting in selective prediction [9, 10]; see Section B for more details. Such a setup is motivated by scenarios such as medical decision-making and knowledge access, where the cost of committing a type-I error (i.e., selecting a non-aligned unit) can be gravy, and should be of primary concern. In addition, we define power as the proportion among truly aligned units, instead of the number of selections—indeed, if a foundation model is not powerful in generating aligned outputs, to begin with, it is natural that only a few of the outputs shall be deployed.
@@ 2.1 Related works
A line of work applies CP to construct prediction sets that cover the outcome for a single test unit in classification or regression problems addressed by language models [23, 25, 43, 45, 51]. While desired, their reliability in downstream tasks is only empirically evaluated: we note that coverage of prediction sets does not carry over to downstream decisions due to selection bias [16, 17]. In contrast, this work applies to outputs of general forms, considers multiple units simultaneously, and offers guarantees relevant to downstream tasks. In particular, our methodology is built upon the series of papers in conformalized selection [16, 15]. While they aim to find large values of outcomes, we motivate and tailor the framework for the alignment of foundation models.
Two works addressing the alignment of foundation models with CP ideas have been briefly discussed in Section 1. Our setup draws ideas from [33], but the guarantees we provide differ from both. In particular, our method may be better suited to situations where it is desirable to avoid modifying the outputs or enlarging candidate sets which requires retraining the models.
@@
Our framework is closely related to the idea of selective prediction, where one is allowed either to predict or to abstain from making a prediction [8, 9, 10, 30, 1]. In the context of large language models, [21, 50, 37, 52] have similarly considered the goal of teaching the model not to predict when the model is uncertain about its output, but theoretical guarantees on the alignment of the selected outputs are lacking. In contrast, this work rigorously formalizes the guarantee and provides an end-to-end framework that achieves it under a mild exchangeability condition. As mentioned earlier, the strict control of FDR is closely connected to the “selection with controlled risk” setting in selective prediction [9, 10], which aims to control a slightly different error measure; we provide a detailed comparison of the error measures and methods in Appendix B.
@@
There is a rapidly growing literature on the alignment of foundation models, wherein topics include alignment evaluation [28, 39], prompt design [20], training with human feedback [2, 32, 35], among others [7]. We rely on commonly-used alignment criteria in our applications. However, we achieve strict error control via post-hoc adjustment/selection, instead of deriving training strategies to improve certain alignment metrics of the model.
@@ 3 Conformal Alignment
The following proposition characterizes the asymptotic power, as well as the fraction of selected units that can be deployed with confidence, when the samples are i.i.d. The proof is in Appendix A.2.
@@
Under the same condition of Theorem 3.1, assume further that $(X_{i},E_{i})_{i\in[n+m]}$ are i.i.d. Define $H(t)=\mathbb{P}(A\leq c,g(X)\geq t)$ and $t(\alpha)=\sup\{t:\frac{t}{\mathbb{P}(H(g(X))\leq t)}\leq\alpha\}$. Suppose that there exists a sufficiently small $\varepsilon>0$ such that $\frac{t}{\mathbb{P}(H(g(X))\leq t)}<q$ for $t\in[t(\alpha)-\varepsilon,t(\alpha)]$, then
@@ 4 Experiments for question answering (QA)
In this section, we implement and evaluate Conformal Alignment in question-answering tasks, where we consider a conversational question answering dataset CoQA [36] and a closed-book reading comprehension dataset TriviaQA [19]. For each QA dataset, we use language models OPT-13B [53] and LLaMA-2-13B-chat [44] without finetuning to generate an answer $f(X_{i})$ via top-p sampling for each input $X_{i}$ following the default configuration. The alignment score $A_{i}$ measures how well the generated answer matches the true reference answers provided in both datasets. Following [22, 27], we let $A_{i}\in\{0,1\}$ indicate whether the rouge-L score [26] between the LLM-generated answer and reference answer is no less than $0.3$, and set $c=0$. See Appendix C.1 for more details.
@@ 4.1 Experiment results
In addition to the capability of LLMs, the features used to train $g$ also play a vital role in the power of selection. To compare the informativeness of the aforementioned features for the alignment of outputs, we use one score at a time as the feature in training $g$ and show Power/FDR curves in Figure 5. When using the OPT-13B model, Self_Eval and Num_Setss are the least powerful, while Deg(J) and Ecc(J) are the most powerful in predicting alignment. Such a comparison is similar for LLaMA-2-13B-chat. We defer the FDR curves to Figure 14 in Appendix D.3 which remain strictly controlled.
@@ 5.1 Experiment results
6 Discussion
We have presented Conformal Alignment, a principled framework for ensuring the alignment of foundation model outputs with provable, distribution-free FDR control. Compared with standard conformal prediction, we provide a solution that may be more relevant to downstream tasks than existing ideas. We demonstrate that our framework is flexible, lightweight, and effective through concrete applications to question answering and radiology report generation tasks, where we also provide practical recommendations on sample size, data splitting, and feature engineering. Interesting future directions include designing implementation tailored for other appropriate applications our framework suits, tackling parallel or sequential outputs, and extending the framework to control other error notions beyond FDR that may be meaningful for downstream decisions.
@@ Acknowledgment
The authors would like to thank the Wharton Research Computing team and Stanford Research Computing Center for the computational resources provided and the great support from the staff members.
Appendix A Deferred proofs
@@
Appendix B Additional comparison with selective prediction

### v3

@@ Abstract
Before deploying outputs from foundation models in high-stakes tasks, it is imperative to ensure that they align with human values. For instance, in radiology report generation, reports generated by a vision-language model must align with human evaluations before their use in medical decision-making. This paper presents Conformal Alignment, [^1] The code is available at https://github.com/yugjerry/conformal-alignment. a general framework for identifying units whose outputs meet a user-specified alignment criterion. It is guaranteed that on average, a prescribed fraction of selected units indeed meet the alignment criterion, regardless of the foundation model or the data distribution. Given any pre-trained model and new units with model-generated outputs, Conformal Alignment leverages a set of reference data with ground-truth alignment status to train an alignment predictor. It then selects new units whose predicted alignment scores surpass a data-dependent threshold, certifying their corresponding outputs as trustworthy. Through applications to question answering and radiology report generation, we demonstrate that our method is able to accurately identify units with trustworthy outputs via lightweight training over a moderate amount of reference data. En route, we investigate the informativeness of various features in alignment prediction and combine them with standard models to construct the alignment predictor.
@@ 1 Introduction
Towards uncertainty quantification of black-box models, conformal prediction [50, CP] offers a versatile, distribution-free solution. While CP has been applied to classification and regression tasks addressed by foundation models [25, 27, 46, 48, 54, 2], its use to ensure the alignment of general-form outputs remains largely unexplored. Recently, [36] propose to generate multiple outputs for one input until it is guaranteed that at least one output meets a specific alignment criterion. Such a guarantee, however, may not be immediately practical, as users must still determine (manually) which output(s) is aligned. [32] modify the machine-generated outputs to ensure factuality with high probability, which in turn may sacrifice model power by making responses vague. [^2] See Appendix D.7 for more detailed discussion on the types of guarantees. As such, determining guarantees that are practically meaningful for downstream tasks and developing methods to achieve them remain an active research area.
This paper introduces a distinct type of guarantee: we aim to certify whether each model output is aligned or not, such that those certified are ensured to be mostly correct. This guarantee is directly relevant to downstream tasks, as it allows immediate and reliable use of the certified outputs without further modifications. On the other hand, we abstain from deploying model outputs for units (data points) that are not selected—intuitively, those are where the model lacks confidence (i.e., “knowing they do not know”), and they can be evaluated by human experts. Such a practice therefore provides an automated pipeline for the safe deployment of model-generated outputs.
Present work.
We present Conformal Alignment (CA), a general framework that determines when to trust foundation model outputs with finite-sample, distribution-free guarantees. The framework is based on CP ideas, but unlike standard CP that searches the sampling space to construct a prediction set that contains a desired output, we leverage the quantified confidence as an instrument to select units whose already-generated outputs are trusted. Given any model and any alignment criterion, Conformal Alignment ensures that a prescribed proportion of the selected units’ model outputs indeed meet the alignment criterion. [^3] In this work, we use “alignment” to refer to a desired property of an output that may vary with the context. Concretely, suppose $m\in\mathbb{N}_{+}$ new units awaiting outputs. Given an error level $\alpha\in(0,1)$, Conformal Alignment controls the false discovery rate (FDR) in selecting units with trustworthy outputs:
@@
We illustrate the pipeline of Conformal Alignment via a use case in radiology report generation (detailed in Section 5) in Figure 1: each prompt corresponds to an X-ray scan, for which a vision-language model generates a report. Our framework identifies a subset of generated reports that are reliable in the sense that at least, say $99\%$, of the selected reports are guaranteed to be aligned if they were to be compared with a human expert report.
@@ 2 Problem setup
For a test set $\mathcal{D}_{\textnormal{test}}=\{X_{n+j}\}_{j=1}^{m}$, we aim to select a subset ${\mathcal{S}}\subseteq[m]:=\{1,\dots,m\}$ such that most of their (unobserved) alignment scores exceed a pre-fixed threshold $c\in\mathbb{R}$. The error is measured by
@@
$\text{Power}=\mathbb{E}\Bigg{[}\frac{\sum_{j\in[m]}\operatorname{\mathds{1}}\{A_{n+j}>c,j\in{\mathcal{S}}\}}{\max(\sum_{j\in[m]}\operatorname{\mathds{1}}\{A_{n+j}>c\},1)}\Bigg{]}.$ (2.1)
It should be emphasized that our framework prioritizes FDR control over power, in that we strictly enforce the former while optimizing the latter under the constraints. This is related to but slightly differs from the selection with controlled risk setting in selective prediction [11, 12]; see Section B.1 for more details. Such a setup is motivated by scenarios such as medical decision-making and knowledge access, where the cost of committing a type-I error (i.e., selecting a non-aligned unit) can be grave, and should be of primary concern. In addition, we define power as the proportion among truly aligned units, instead of the number of selections—indeed, if a foundation model is not powerful in generating aligned outputs, to begin with, it is natural that only a few of the outputs shall be deployed.
@@ 2.1 Related works
A line of work applies CP to construct prediction sets that cover the outcome for a single test unit in classification or regression problems addressed by language models [25, 27, 46, 48, 54]. While desired, their reliability in downstream tasks is only empirically evaluated: we note that coverage of prediction sets does not carry over to the selected units [18, 19]. In contrast, this work applies to outputs of general forms, considers multiple units simultaneously, and offers guarantees relevant to downstream tasks. In particular, our methodology is built upon the series of papers in conformalized selection [18, 17]. While they aim to find large values of outcomes, we motivate and tailor the framework for the alignment of foundation models (a more detailed discussion is in Appendix B.2).
Two works [36, 32] addressing the alignment of foundation models with CP ideas have been briefly discussed in Section 1. Our setup draws ideas from [36], but the guarantees we provide differ from both. In particular, our method may be better suited to situations where it is desirable to avoid modifying the outputs or enlarging candidate sets which requires retraining the models.
@@
Our framework is closely related to the idea of selective prediction, where one is allowed either to predict or to abstain from making a prediction [9, 11, 12, 33, 1]. In the context of large language models, [23, 53, 40, 55] have similarly considered the goal of teaching the model not to predict when the model is uncertain about its output, but theoretical guarantees on the alignment of the selected outputs are lacking. In contrast, this work rigorously formalizes the guarantee and provides an end-to-end framework that achieves it under a mild exchangeability condition. As mentioned earlier, the strict control of FDR is closely connected to the “selection with controlled risk” setting in selective prediction [11, 12], which aims to control a slightly different error measure; we provide a detailed comparison of the error measures and methods in Appendix B.1.
@@
There is a rapidly growing literature on the alignment of foundation models, wherein topics include alignment evaluation [30, 42], prompt design [22], training with human feedback [3, 35, 38], among others [8]. We rely on commonly-used alignment criteria in our applications [44, 29, 24]. However, we achieve strict error control via post-hoc adjustment/selection, instead of deriving training strategies to improve certain alignment metrics of the model.
@@ 3 Conformal Alignment
The following proposition characterizes the power (2.1), as well as the fraction of selected units that can be deployed with confidence, when the samples are i.i.d. The proof is in Appendix A.2. A finite-sample power analysis based on concentration inequalities is in Appendix A.3.
@@
Under the same condition of Theorem 3.1, assume further that $(X_{i},E_{i})_{i\in[n+m]}$ are i.i.d. Define $H(t)=\mathbb{P}(A\leq c,g(X)\geq t)$ and $t(\alpha)=\sup\{t:\frac{t}{\mathbb{P}(H(g(X))\leq t)}\leq\alpha\}$. Suppose that there exists a sufficiently small $\varepsilon>0$ such that $\frac{t}{\mathbb{P}(H(g(X))\leq t)}<\alpha$ for $t\in[t(\alpha)-\varepsilon,t(\alpha)]$, then
@@ 4 Experiments for question answering (QA)
In this section, we implement and evaluate Conformal Alignment in question-answering tasks, where we consider a conversational question answering dataset TriviaQA [21] and a closed-book reading comprehension dataset CoQA [39]. For each QA dataset, we use language models OPT-13B [57] and LLaMA-2-13B-chat [47] without finetuning to generate an answer $f(X_{i})$ via top-p sampling for each input $X_{i}$ following the default configuration. The alignment score $A_{i}$ measures how well the generated answer matches the true reference answers provided in both datasets. Following [24, 29], we let $A_{i}\in\{0,1\}$ indicate whether the rouge-L score [28] between the LLM-generated answer and reference answer is no less than $0.3$, and set $c=0$. See Appendix C.1 for more details.
@@ 4.1 Experiment results
In addition to the capability of LLMs, the features used to train $g$ also play a vital role in the power of selection. To compare the informativeness of the aforementioned features for the alignment of outputs, we use one score at a time as the feature in training $g$ and show Power/FDR curves in Figure 5. When using the OPT-13B model, Self_Eval and Num_Setss are the least powerful, while Deg(J) and Ecc(J) are the most powerful in predicting alignment. Such a comparison is similar for LLaMA-2-13B-chat. We defer the FDR curves to Figure 14 in Appendix D.3 which remain strictly controlled. We include a more detailed analysis on feature importance via Shapley value in Appendix D.6.
@@ 5.1 Experiment results
6 Discussion and limitations
We have presented Conformal Alignment, a principled framework for ensuring the alignment of foundation model outputs with provable, distribution-free FDR control. Compared with standard conformal prediction, we provide a solution that may be more relevant to downstream tasks than existing ideas. We demonstrate that our framework is flexible, lightweight, and effective through concrete applications to question answering and radiology report generation tasks, where we also provide practical recommendations on sample size, data splitting, and feature engineering.
The choice of alignment score in different applications remains an open problem. For example, in the radiology report generation example, as pointed out by [56], there can be a discrepancy between the score given by CheXbert and that given by human evaluation; as such, it will be interesting to use our framework with actual human evaluation labels. More generally, interesting future directions include designing implementation tailored for other appropriate applications, tackling parallel or sequential outputs, improving power in data-scarce scenarios (e.g., with better base models and more informative features), extending to multi-task and in-context learning settings, and generalizing our framework to control other error notions beyond FDR that may be meaningful for downstream decisions.
@@ Acknowledgment
Z.R. is supported by NSF grant DMS-2413135. The authors would like to thank the Wharton Research Computing team and Stanford Research Computing Center for the computational resources provided and the great support from the staff members.
Appendix A Deferred theory and proofs
@@
A.3 Finite-sample power analysis
In this part, we provide a finite-sample power analysis based on concentration inequalities. We define
$F_{+}(t)=\mathbb{P}(g(X)\geq t,A>c),\quad F_{-}(t)=\mathbb{P}(g(X)\geq t,A\leq c),\quad F(t)=\mathbb{P}(g(X)\geq t).$
The empirical version of these quantities are defined as
$\widehat{F}_{m}^{+}(t)=\frac{1}{m}\sum_{j=1}^{m}\operatorname{\mathds{1}}\{g(X_{n+j})\geq t,A_{n+j}>c),\quad\widehat{F}_{n}^{+}(t)=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathds{1}}\{g(X_{i})\geq t,A_{i}>c),$
$\widehat{F}_{m}^{-}(t)=\frac{1}{m}\sum_{j=1}^{m}\operatorname{\mathds{1}}\{g(X_{n+j})\geq t,A_{n+j}\leq c),\quad\widehat{F}_{n}^{-}(t)=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathds{1}}\{g(X_{i})\geq t,A_{i}\leq c),$
$\widehat{F}_{m}(t)=\frac{1}{m}\sum_{j=1}^{m}\operatorname{\mathds{1}}\{g(X_{n+j})\geq t),\quad\widehat{F}_{n}(t)=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathds{1}}\{g(X_{i})\geq t),$
We also define $\widehat{m}^{+}=\sum_{j=1}^{m}\operatorname{\mathds{1}}\{A_{n+j}>c\}$ as the number of truly aligned test outputs.
Proposition A.1.
Under the same condition of Theorem 3.1, assume further that $(X_{i},E_{i})_{i\in[n+m]}$ are i.i.d. For any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\frac{F_{+}(\tau_{-})-\epsilon_{m}}{\mathbb{P}(A>c)+\epsilon_{m}}\leq\frac{\sum_{j=1}^{m}\operatorname{\mathds{1}}\{g(X_{n+j})\geq\widehat{\tau},A_{n+j}>c\}}{\sum^{m}_{j=1}\operatorname{\mathds{1}}\{A_{n+j}>c\}}\leq\frac{F_{+}(\tau_{+})+\epsilon_{m}}{\mathbb{P}(A>c)-\epsilon_{m}},$
where
$\tau_{+}:=\inf\bigg{\{}t\colon\frac{1}{n+1}\cdot\frac{1+nF_{+}(t)+n\epsilon_{n}}{F_{m}(t)-\epsilon_{m}}\leq\alpha\bigg{\}},\quad\tau_{-}:=\inf\bigg{\{}t\colon\frac{1}{n+1}\cdot\frac{1+nF_{+}(t)-n\epsilon_{n}}{F_{m}(t)+\epsilon_{m}}\leq\alpha\bigg{\}},$
and $\epsilon_{m}=\sqrt{\log(8/\delta/(2n))}$, $\epsilon_{n}=\sqrt{\log(8/\delta)/(2m)}$.
Proof of Proposition A.1.
Fix any constant $\delta\in(0,1)$. Consider the function $\widetilde{g}(X,A)=M\cdot\operatorname{\mathds{1}}\{A>c\}+g(X)$, where $M>2\sup_{x\in\mathcal{X}}|g(x)|<\infty$ is a sufficiently large constant such that $\inf_{x\in\mathcal{X}}\widetilde{g}(x,c+1)>\sup_{x\in\mathcal{X}}\widetilde{g}(x,c)$. Therefore, we know that for any $|t|<M/2$, it holds that
$F_{+}(t)=\mathbb{P}(\widetilde{g}(X,A)\geq M+t),\quad F_{-}(t)=\mathbb{P}(t\leq\widetilde{g}(X,A)<M/2).$
And similar relations hold for the empirical versions of these quantities. Applying the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality [10, 31] with a union bound for the empirical c.d.f. of $\widetilde{g}(X,A)$ for both calibration and test data, we know that with probability at least $1-\delta$,
$\sup_{t}\big{|}\widehat{F}_{m}^{+}(t)-F_{+}(t)\big{|}\leq\epsilon_{m},\quad\sup_{t}\big{|}\widehat{F}_{m}^{-}(t)-F_{-}(t)\big{|}\leq\epsilon_{m},\quad\sup_{t}\big{|}\widehat{F}_{m}(t)-F(t)\big{|}\leq\epsilon_{m},$
$\sup_{t}\big{|}\widehat{F}_{n}^{+}(t)-F_{+}(t)\big{|}\leq\epsilon_{n},\quad\sup_{t}\big{|}\widehat{F}_{n}^{-}(t)-F_{-}(t)\big{|}\leq\epsilon_{n},\quad\textrm{and}\quad\big{|}\widehat{m}^{+}-\mathbb{P}(A>c)\big{|}\leq\epsilon_{m},$
where recall that $\epsilon_{m}=\sqrt{\log(8/\delta)/(2m)}$ and $\epsilon_{n}=\sqrt{\log(8/\delta)/(2n)}$. We call the event that all above holds the “good event” $E^{*}$, which happens with probability at least $1-\delta$.
By definition, we can show that the selection set of Conformal Alignment is equivalently ${\mathcal{S}}=\{j\in[m]\colon g(X_{n+j})\geq\widehat{\tau}\}$, where $\widehat{\tau}=\inf\big{\{}t\colon\widehat{\text{FDR}}(t)\leq\alpha\big{\}}$, and
$\widehat{\text{FDR}}(t)=\frac{m}{n+1}\cdot\frac{1+\sum_{i=1}^{n}\operatorname{\mathds{1}}\{g(X_{i})\geq t,A_{i}\leq c\}}{\sum_{j=1}^{m}\operatorname{\mathds{1}}\{g(X_{n+j})\geq t\}}=\frac{1}{n+1}\cdot\frac{1+n\cdot\widehat{F}_{n}^{+}(t)}{\widehat{F}_{m}(t)}.$
By the DKW inequality above, with probability at least $1-\delta$,
$\frac{1}{n+1}\cdot\frac{1+nF_{+}(t)-n\epsilon_{n}}{F_{m}(t)+\epsilon_{m}}\leq\widehat{\text{FDR}}(t)\leq\frac{1}{n+1}\cdot\frac{1+nF_{+}(t)+n\epsilon_{n}}{F_{m}(t)-\epsilon_{m}}$
holds simultaneously for all $t\in\mathbb{R}$. On the other hand, the (realized) power is
$\text{Pow}:=\frac{\sum_{j=1}^{m}\operatorname{\mathds{1}}\{g(X_{n+j})\geq\widehat{\tau},A_{n+j}>c\}}{\sum_{j=1}^{m}\operatorname{\mathds{1}}\{A_{n+j}>c\}}=\frac{\widehat{F}_{m}^{+}(\widehat{\tau})}{m^{+}}.$
On the good event $E^{*}$, we have
$\frac{F_{+}(\widehat{\tau})-\epsilon_{m}}{\mathbb{P}(A>c)+\epsilon_{m}}\leq\text{Pow}\leq\frac{F_{+}(\widehat{\tau})+\epsilon_{m}}{\mathbb{P}(A>c)-\epsilon_{m}}.$
By the definition of $\widehat{\tau}$, we know that $\tau_{-}\leq\widehat{\tau}\leq\tau_{+}$ on the good event $E^{*}$. The monotonicity of realized power as a function of selection cutoff $\widehat{\tau}$ thus implies that
$\text{Pow}\geq\frac{\widehat{F}_{m}^{+}(\tau_{-})}{m^{+}}\geq\frac{F_{+}(\tau_{-})-\epsilon_{m}}{\mathbb{P}(A>c)+\epsilon_{m}},$
$\text{Pow}\leq\frac{\widehat{F}_{m}^{+}(\tau_{+})}{m^{+}}\leq\frac{F_{+}(\tau_{+})+\epsilon_{m}}{\mathbb{P}(A>c)-\epsilon_{m}}.$
Appendix B Additional comparison with prior works
B.1 Comparison with selective prediction